# references.bib

### v5

@@ 1 Introduction
Price impact refers to the fact that buyers push the price up and sellers push the price down [bouchaud2010impact, bouchaud2018trades]. The traditional Efficient Market interpretation of this empirical fact is that buyers and sellers are on average informed, and lo and behold, the price moves according to their prediction [hasbrouck1991measuring, hasbrouck2007empirical]. This scenario is at the heart of the celebrated Kyle model for price impact [kyle1985continuous].
Another, very different interpretation of price impact is that it is a purely statistical reaction of the market to incoming order flow, where information plays little role. Prices move simply because people trade, whatever the reason they are trading, and volatility is the result of people randomly buying and selling. This is the Order-Driven view of markets, explicitly spelled out in [bouchaud2018trades], chapter 20 – but see also [bouchaud2006random, hopman2007supply, bouchaud2010impact] and in a different setting, [gabaix2021search, bouchaud2022inelastic]. In this scenario, there is no “information revelation” but rather “self-fulfilling prophecies”, as emphasized in [bouchaud2009markets, van2024ponzi].
Of course, reality should lie somewhere between these two extremes. There are surely some informed trades, and news do impact prices, but there is also overwhelming evidence for the presence of “noise traders”, excess trading and excess volatility in financial markets, see e.g. [black1986noise, odean1999investors, shiller1981stock, leroy2006excess, kurth2025revisiting]. Direct empirical estimates suggest that informed trades are a minority (see [bouchaud2003fluctuations] and [bouchaud2018trades], chapter 16). While we are convinced that the Order-Driven view is a much closer approximation to the dynamics of markets, there are several empirical loose ends that need to be tied up. A major issue is how to reconcile two prominent stylized facts of market microstructure, namely, (i) the long-range memory in order signs (i.e. $+1$ for buy orders and $-1$ for sell orders) and (ii) the ubiquitous square-root law of market impact (that governs the average price move induced by the execution of a sequence of orders) with the random walk nature of prices, with a volatility $\sigma$ that is directly proportional to the amplitude of the square-root law. More precisely, the square-root law states that the average price impact $\mathcal{I}$ of a metaorder of total volume $Q$ is given by
@@
where $Y$ is a $O(1)$ numerical coefficient and $\phi$ is the average flow of orders executed in the market per unit time – see e.g. [loeb1983trading, grinold2000active, almgren2005direct, toth2011anomalous, bacry2015market, donier2015million, toth2016square, bucci2018slow, AQR, sato2024does, maitrier2025generatingrealisticmetaorderspublic] and [bouchaud2018trades, webster2023handbook] for reviews.
Eq. (1) may look familiar, and in fact trivial: superficially, it states that price changes (i.e. $\mathcal{I}(Q)$) grow as the square-root of execution time $T$, i.e. the law of random walks, provided one assumes that $T$ and $Q$ are proportional. But, as argued in [bucci_impact_vol2019], such an argument is completely misleading: not only $\mathcal{I}(Q)$ is an average price change and not a standard deviation; but also $\mathcal{I}(Q)$ is found to depend only on the quantity executed $Q$ and not on execution time $T$. Furthermore, it is known that the impact of metaorders decays, on average, after the end of the execution period (e.g. [moro2009market, brokmann2015slow, bucci2018slow]). Note that all these features are at odds with the Kyle model that predicts linear and permanent impact, resulting from information revelation [kyle1985continuous]. [^1] Related to this discussion, see also [saddier2024bayesian].
@@
Although many theories have been proposed to explain the square-root law, there is no consensus on the issue. The Latent Liquidity Theory proposed in [donier2015fullyconsistentminimalmodel] (see also [donier2016walras, bouchaud2018trades]) seems to capture many features observed in data but seems inconsistent with others, in particular those reported in the recent paper of Maitrier et al. [maitrier2025double]. We will not attempt to dwell further on this particular issue here but accept it as an incontrovertible empirical fact, still waiting for a fully convincing explanation. We will rather focus on points (b) and (c) above: knowing that the duration of metaorders is power-law distributed, can one reconcile the square-root impact law with the volatility of markets and with a locally linear aggregate impact law?
In order to answer these questions quantitatively, we introduce in section 2 a new theoretical framework to describe metaorders with different signs, sizes and durations, which impact prices as a square-root of volume but with a subsequent time decay. We show in section 4 that, as in the original propagator model, price diffusion is ensured by the long memory of cross-correlations between metaorders. In order to account for the effect of strongly fluctuating volumes $q$ of individual trades, we need to further introduce two $q$-dependent exponents, which we justify empirically and allow us to account for the way the moments of generalized volume imbalance (section 3) and the correlation between price changes and generalized volume imbalance (section 5) scales with $T$. We predict in particular that the corresponding power-laws depend in a non-monotonic fashion on a parameter $a$ that allows one to put the same weight on all child orders or overweight large orders, a behaviour clearly borne out by empirical data (section 5). We also predict that the correlation between price changes and volume imbalances should display a maximum as a function of $a$, which again matches observations (section 5). We conclude by arguing that our results support the “Order-Driven” theory of markets, and are at odds with the idea that a “Fundamental” component accounts for a large share of the volatility of financial markets.
@@ 2.1 Model set-up
We posit that between $t$ and $t+{\rm d}t$ and with probability $\nu{\rm d}t$ a new metaorder of random sign $\varepsilon(t)=\pm 1$ and duration $s(t)$ is initiated. The volume of child orders is $q$ (which might itself be random, see below), and during execution the probability that one of them gets executed is $\varphi{\rm d}t$, independently of the size $q$. We neglect throughout this paper activity fluctuations as well as intraday seasonalities, as these are not crucial for the effects we want to focus on. This means that $\mu,\varphi$ and the average duration $\bar{s}$ are chosen to be time independent.
The total size of the metaorder is thus $Q=q\varphi s+O(\sqrt{s})$. The probability density of durations $s$ is denoted $\Psi(s)$, which will typically has a power-law tail $\Psi(s)\propto s^{-1-\mu}$, such that the distribution of metaorder sizes $Q$ inherits from this power-law, and decays as $Q^{-1-\mu}$ as suggested by empirical data [farmer2013efficiency, sato2023inferring].
Such a power-law distribution of metaorder sizes is the basic mechanism proposed by Lillo, Mike and Farmer (LMF) [lillo2005theory] to explain the long memory of order signs, which is known to decay with lag $\tau$ as $\tau^{-\gamma}$ with $10<\gamma<$ 1 [bouchaud2009markets]. Within the LMF model, one has $\gamma=\mu-1$, a result recently validated in great details by Sato and Kanazawa [sato2023inferring] using data from the Tokyo Stock Exchange. In a later stage, we will allow the exponent $\mu$ to depend on $q$, to account for the fact that large child orders tend to be less autocorrelated that small ones.
We will also allow the sign of different metaorders to be correlated, as indeed observed in data [donier2015million, bucci2020co]. More precisely, we will model the long-term decay of the autocorrelation of signs, $\mathbb{E}[\varepsilon(t)\varepsilon(t+\tau)]$ as a power-law $\tau^{-\gamma_{\times}}$, with an exponent $\gamma_{\times}$ a priori such that $\gamma_{\times}\geq\gamma$ such not to contradict the LMF hypothesis.
@@ 2.2 Average number of metaorders
In the following, we will always make averages over the metaorder initiation process, and often replace ${\rm d}N_{t}$ by $\nu{\rm d}t$ whenever possible. When comparing with empirical data, we will work in trade time $N_{T}$ but still call this quantity $T$. Translating our results is real time is, however, non-trivial because, as is well known, the activity rate $\nu$ shows strongly intermittent dynamics (often modeled using Hawkes processes, see e.g. [bouchaud2018trades], chapter 9) on top of a U-shaped intraday pattern.
@@ 2.3 Average trading activity and trading volume
Carefully taking the derivative with respect to $T$ one finds:
$\frac{\partial\overline{V}_{T}^{1}}{\partial T}=\kappa\left[\int_{0}^{T}{\rm d}s\,s\Psi(s)+T\int_{T}^{\infty}{\rm d}s\,\Psi(s)\right];\qquad\frac{\partial\overline{V}_{T}^{2}}{\partial T}=\kappa\int_{T}^{\infty}{\rm d}s\,(s-T)\Psi(s).$ (9)
Hence, the total average executed volume $\overline{V}_{T}=\overline{V}_{T}^{1}+\overline{V}_{T}^{2}$ is given, for large $T$, by
@@ 3.1 Sign Imbalance
corresponding to a sign autocorrelation function decaying as $\tau^{-\gamma}$ with $\gamma=\mu-1<1$, as found in the data [lillo2005theory, sato2023inferring]. Then the previous expression becomes:
@@
When $\mu<4$, this behaves as $T^{4-\mu}$, so that the kurtosis of the sign imbalance distribution behaves as $T^{5-\mu}/(T^{3-\mu})^{2}\sim T^{\mu-1}$ which grows with $T$! An important consequence is that within our model the sign imbalance does not become Gaussian for large $T$.
Generalizing to the $2n$-th moment, one finds that it grows with $T$ like $T^{2n+1-\mu}$ when $\mu<2n$. This suggests that when $\mu<2$, the sign imbalance converges at large $T$ towards a truncated Lévy distribution of index $\mu$ for the rescaled variable $I^{0}/T^{1/\mu}$, where the truncation takes place for $|I^{0}|=T$ (see [Wyart_2003] for a very similar calculation, and the Appendix of [moran2024revisitinggranularmodelsfirm] for a proof). Indeed, one can check that the moments of such a truncated Lévy distribution scale with $T$ exactly as above. We will test this prediction in section 3.4.
@@ 3.2 Generalized Volume Imbalance
Note that $a=0$ corresponds to sign imbalances and $a=1$ to volume imbalances. Obviously, if $q(t)=q$ at all times, all these imbalances are equal, up to a trivial factor, to the sign imbalance computed in the previous section. In the following, we will assume that metaorders differ not only by their duration $s$ but also by the size of their child orders, with a joint distribution that we denote as $\Psi_{q}(s)\Xi(q)$. Inspired by empirical data (see below), we posit that metaorders that execute with larger child order sizes $q$ still have a power-law distributed duration $s$, but with a tail exponent $\mu_{q}$ that increases with $q$ – i.e. have a thinner tail. More precisely, the conditional distribution $\Psi_{q}(s)$ is of the form:
$\Psi_{q}(s)=\frac{\mu_{q}s_{0}^{\mu_{q}}}{s^{1+\mu_{q}}},\qquad\mu_{q}=\mu_{1}+\lambda\log q,$ (22)
where $q=1$ is the lot size.
@@
[…]

### v6

# Order Imbalance & Volatility: A Unifying Framework

@@ Abstract
Keywords: Market microstructure, price impact, square-root law, order flow dynamics, diffusivity puzzle, physics of financial markets.
@@ 1 Introduction
Price impact refers to the fact that buyers push the price up and sellers push the price down [9, 5]. The traditional Efficient Market interpretation of this empirical fact is that buyers and sellers are on average informed, and lo and behold, the price moves according to their prediction [27, 28]. This scenario is at the heart of the celebrated Kyle model for price impact [32].
Another, very different interpretation of price impact is that it is a purely statistical reaction of the market to incoming order flow, where information plays little role. Prices move simply because people trade, whatever the reason they are trading, and volatility is the result of people randomly buying and selling. This is the Order-Driven view of markets, explicitly spelled out in [5], chapter 20 – but see also [8, 29, 9] and in a different setting, [24, 10]. In this scenario, there is no “information revelation” but rather “self-fulfilling prophecies”, as emphasized in [6, 53].
Of course, reality should lie somewhere between these two extremes. There are surely some informed trades, and news do impact prices, but there is also overwhelming evidence for the presence of “noise traders”, excess trading and excess volatility in financial markets, see e.g. [4, 41, 48, 33, 31]. Direct empirical estimates suggest that informed trades are a minority (see [7] and [5], chapter 16). While we are convinced that the Order-Driven view is a much closer approximation to the dynamics of markets, there are several empirical loose ends that need to be tied up. A major issue is how to reconcile two prominent stylized facts of market microstructure, namely, (i) the long-range memory in order signs (i.e. +1+1 for buy orders and 1-1 for sell orders) and (ii) the ubiquitous square-root law of market impact (that governs the average price move induced by the execution of a sequence of orders) with the random walk nature of prices, with a volatility $\sigma$ that is directly proportional to the amplitude of the square-root law. More precisely, the square-root law states that the average price impact $\mathcal{I}$ of a metaorder of total volume $Q$ is given by
@@
where $Y$ is a $O(1)$ numerical coefficient and $\phi$ is the average flow of orders executed in the market per unit time – see e.g. [35, 26, 1, 51, 3, 18, 50, 12, 23, 46, 36] and [5, 54] for reviews.
Eq. (1) may look familiar, and in fact trivial: superficially, it states that price changes (i.e. $\mathcal{I}(Q)$) grow as the square-root of execution time $T$, i.e. the law of random walks, provided one assumes that $T$ and $Q$ are proportional. But, as argued in [14], such an argument is completely misleading: not only $\mathcal{I}(Q)$ is an average price change and not a standard deviation; but also $\mathcal{I}(Q)$ is found to depend only on the quantity executed $Q$ and not on execution time $T$. Furthermore, it is known that the impact of metaorders decays, on average, after the end of the execution period (e.g. [39, 11, 12]). Note that all these features are at odds with the Kyle model that predicts linear and permanent impact, resulting from information revelation [32]. [^1] Related to this discussion, see also [44].
@@
Although many theories have been proposed to explain the square-root law, there is no consensus on the issue. The Latent Liquidity Theory proposed in [17] (see also [19, 5]) seems to capture many features observed in data but seems inconsistent with others, in particular those reported in the recent paper of Maitrier et al. [37] using the Tokyo Stock Exchange detailed database. We will not attempt to dwell further on this particular issue here but accept it as an incontrovertible empirical fact, still waiting for a fully convincing explanation. We will rather focus on points (b) and (c) above: knowing that the duration of metaorders is power-law distributed, can one reconcile the square-root impact law with the volatility of markets and with a locally linear aggregate impact law?
2 Main contributions, key results and roadmap
This paper proposes a unified, metaorder-based framework that connects four empirical pillars of market microstructure: (i) the square-root law of metaorder impact, (ii) the post-execution decay of impact, (iii) the long memory of order flow, and (iv) diffusive prices. The objective is not to derive the square-root law from first principles, but to study its consequences when combined with heavy-tailed metaorder durations, heterogeneous trade sizes, and (crucially) correlations between metaorders.
Main results. The technical developments of the paper lead to three takeaways:
  1. 1.
    Diffusion with transient square-root impact requires cross-metaorder correlations. When metaorder impact is transient (decaying after execution), the heavy-tailed distribution of metaorder durations alone leads to mean-reversion, i.e. subdiffusive impact-induced variance (Section 5). We show that price diffusion is restored when the signs of different metaorders are long-range correlated, $\mathbb{E}[\varepsilon(t)\varepsilon(t+\tau)]\sim\tau^{-\gamma_{\times}}$, because the resulting contribution to price variance scales as $T^{2-\gamma_{\times}-2\beta}$ (Eq. (55)). This yields diffusion for $\gamma_{\times}\simeq 1-2\beta$, which is empirically falsifiable using, e.g. the Tokyo Stock exchange data of [45, 46].
  2. 2.
    Volume heterogeneity induces non-trivial, $a$-dependent imbalance scaling. To bridge the gap between sign $\varepsilon$ imbalance and volume $q$ imbalance, we introduce the generalized imbalance $I_{T}^{a}=\sum_{t\in[0,T]}\varepsilon_{t}(q_{t})^{a}$ (with exponent $a=0$ and $a=1$ as special cases). Motivated by data, we allow the metaorder-duration tail exponent to depend on the typical child order size, $\mu_{q}$, in such a way that large-$q$ orders are empirically less persistent. This mechanism predicts an $a$-dependent crossover in the scaling of the moments of $I_{T}^{a}$ (Eqs. (31)–(33)), confirmed empirically (Fig. 3).
  3. 3.
    Reconciling square-root metaorder impact with locally linear aggregated impact observables. The above ingredients yield testable predictions for covariances and correlations between returns $\Delta_{T}$ and generalized imbalances $I_{T}^{a}$. In particular, the effective scaling exponent of $\mathbb{E}[\Delta_{T}\cdot I_{T}^{a}]$ as a function of $T$ is predicted to be non-monotonic in $a$ (Section 6, Eqs. (66)–(67)), and the correlation coefficient $R_{a}(T)$ is predicted to be hump-shaped as a function of $a$ (Section 6.3, Eq. (74)). Both features are observed in the data (Figs. 79).
Relation to the “double square-root law”. Recent evidence suggests that (i) peak impact follows a square-root in executed volume and (ii) the instantaneous impact of a child order scales as $\sqrt{q}$ [37]. Our framework is designed to be consistent with this empirical input: we take $\theta(q)\propto\sqrt{q}$ (Sections 55.3), and study how such a microscopic scaling propagates to mesoscopic observables (diffusion, covariances, correlations) once metaorders overlap and relax.
Implications for volatility generation (order-driven vs. fundamental). Because diffusion can be obtained from the superposition of (possibly correlated) metaorders with transient impact, the model naturally predicts an impact amplitude proportional to volatility (Eq. (57)). We also derive how an additional exogenous “fundamental” component would affect covariances/correlations (Section 6.2), yielding scalings that are not supported by the observed $a$-dependence. This provides empirical evidence that a substantial fraction of volatility is mechanically generated by trading activity.
Roadmap
To address questions (b) and (c) quantitatively, we first introduce in Section 3 a continuous-time description of the order flow as an overlapping population of metaorders with heterogeneous signs, durations and child-order sizes. Throughout, we take as an empirical input that each metaorder generates a square-root peak impact in executed volume, followed by a decay after execution.
In Section 4, we study the statistics of (generalized) order-flow imbalances over windows of size $T$. In particular, to account for the strong heterogeneity of trade sizes $q$, we introduce a family of imbalances $I_{T}^{a}$ that interpolates between sign imbalance ($a=0$) and volume imbalance ($a=1$). This leads us to introduce two $q$-dependent exponents (empirically motivated and estimated from the data) that govern both the persistence of order flow and the decay of impact. These ingredients yield non-trivial scaling predictions for the moments of $I_{T}^{a}$ as a function of $T$.
In Section 5, we then turn to the impact–diffusivity puzzle: how can transient, concave (square-root) impact be compatible with diffusive prices? We show that, in close analogy with the original propagator mechanism, diffusion is restored by the long memory of cross-correlations between distinct metaorders. This identifies metaorder cross-correlations as a crucial ingredient of an order-driven volatility mechanism when impact decays.
Finally, in Sections 66.3, we connect the above building blocks to observables that can be measured directly from trade and price data, namely covariances and correlations between returns and generalized imbalances. We predict that the relevant scaling exponents depend non-monotonically on the weighting parameter $a$ (which interpolates between equal weight per trade and overweighting large trades), and that the correlation between returns and imbalances is hump-shaped as a function of $a$. Both predictions are borne out empirically.
We conclude by discussing the implications of these results for volatility generation: the observed scalings and $a$-dependences are naturally explained when volatility is predominantly impact-driven, and are hard to reconcile with the hypothesis that an exogenous “fundamental” component accounts for a large fraction of price fluctuations. A simulation-based companion paper [maitrier2025subtle] broadly confirms the theoretical results presented here and implements the framework as an artificial market generator.
@@ 3.1 Model set-up
We posit that between $t$ and $t+{\rm d}t$ and with probability $\nu{\rm d}t$ a new metaorder of random sign $\varepsilon(t)=\pm 1$ and duration $s(t)$ is initiated. The volume of child orders is $q$ (which might itself be random, see below), and during execution the probability that one of them gets executed is $\varphi{\rm d}t$, independently of the size $q$.
The total size of the metaorder is thus $Q=q\varphi s+O(\sqrt{s})$. The probability density of durations $s$ is denoted $\Psi(s)$, which will typically has a power-law tail $\Psi(s)\propto s^{-1-\mu}$, such that the distribution of metaorder sizes $Q$ inherits from this power-law, and decays as $Q^{-1-\mu}$ as suggested by empirical data [22, 45]. We neglect throughout this paper activity fluctuations as well as intraday seasonalities, as these are not crucial for the effects we want to focus on. This means that $\mu,\varphi$ and the average duration $\bar{s}$ are chosen to be time independent.
Such a power-law distribution of metaorder sizes is the basic mechanism proposed by Lillo, Mike and Farmer (LMF) [34] to explain the long memory of order signs, which is known to decay with lag $\tau$ as $\tau^{-\gamma}$ with $0<\gamma<1$ [6]. Within the LMF model, one has $\gamma=\mu-1$, a result recently validated in great details by Sato and Kanazawa [45] using data from the Tokyo Stock Exchange. In a later stage, we will allow the exponent $\mu$ to depend on $q$, to account for the fact that large child orders tend to be less autocorrelated that small ones.
We will also allow the sign of different metaorders to be correlated, as indeed observed in data [18, 15]. More precisely, we will model the long-term decay of the autocorrelation of signs, $\mathbb{E}[\varepsilon(t)\varepsilon(t+\tau)]$ as a power-law $\tau^{-\gamma_{\times}}$, with an exponent $\gamma_{\times}$ a priori such that $\gamma_{\times}\geq\gamma$ such not to contradict the LMF hypothesis.
@@ 3.2 Average number of metaorders
In the following, we will always make averages over the metaorder initiation process, and often replace ${\rm d}N_{t}$ by $\nu{\rm d}t$ whenever possible. When comparing with empirical data, we will work in trade time $N_{T}$ but still call this quantity $T$. Translating our results is real time is, however, non-trivial because, as is well known, the activity rate $\nu$ shows strongly intermittent dynamics (often modeled using Hawkes processes, see e.g. [5], chapter 9) on top of a U-shaped intraday pattern.
@@ 3.3 Average trading activity and trading volume
Carefully taking the derivative with respect to $T$ one finds that the total average executed volume $\overline{V}_{T}=\overline{V}_{T}^{1}+\overline{V}_{T}^{2}$ is given, for large $T$, by
@@ 4.1 Sign Imbalance
corresponding to a sign autocorrelation function decaying as $\tau^{-\gamma}$ with $\gamma=\mu-1<1$, as found in the data [34, 45]. Then the previous expression becomes:
@@
When $\mu<4$, this behaves as $T^{4-\mu}$, so that the kurtosis of the sign imbalance distribution behaves as $T^{5-\mu}/(T^{3-\mu})^{2}\sim T^{\mu-1}$ which grows with TT! An important consequence is that within our model the sign imbalance does not become Gaussian for large $T$.
Generalizing to the $2n$-th moment, one finds that it grows with $T$ like $T^{2n+1-\mu}$ when $\mu<2n$. This suggests that when $\mu<2$, the sign imbalance converges at large $T$ towards a truncated Lévy distribution of index $\mu$ for the rescaled variable $I^{0}/T^{1/\mu}$, where the truncation takes place for $|I^{0}|=T$ (see [56] for a very similar calculation, and the Appendix of [38] for a proof). Indeed, one can check that the moments of such a truncated Lévy distribution scale with $T$ exactly as above. We will test this prediction in section 4.4.
@@ 4.2 Generalized Volume Imbalance
Note that $a=0$ corresponds to sign imbalances and $a=1$ to volume imbalances. Obviously, if $q(t)=q$ at all times, all these imbalances are equal, up to a trivial factor, to the sign imbalance computed in the previous section. In the following, we will assume that metaorders differ not only by their duration $s$ but also by the size of their child orders, with a joint distribution that we denote as $\Psi_{q}(s)\Xi(q)$. Inspired by empirical data, we posit that metaorders that execute with larger child order sizes $q$ still have a power-law distributed duration $s$, but with a tail exponent $\mu_{q}$ that increases with $q$ – i.e. have a thinner tail. More precisely, the conditional distribution $\Psi_{q}(s)$ is of the form:
$\Psi_{q}(s)=\frac{\mu_{q}s_{0}^{\mu_{q}}}{s^{1+\mu_{q}}},\qquad\mu_{q}=\mu_{1}+\lambda\ell,$ (21)
where $\ell=\log q$ and $q=1$ corresponds to the lot size.  [^2] Intuitively, it makes sense the sign of large child orders should have shorter memory than small ones. Traders who have large quantities $Q$ to execute is likely to trade small lots in order not to reveal information, whereas smaller $Q$ might be possible to execute in a few shots.
@@
In other words, one finds that the variance of the generalized volume imbalance scales anomalously with $T$ when $a$ is small enough (like the sign imbalance considered above), but becomes simply diffusive when $a$ is large. Intuitively, it is because large child orders are much less auto-correlated than small child orders when $\lambda>0$. [^4] Another mechanism that leads to dependence of the effective exponent of $\Sigma_{a}^{2}(T)$ is the presence of power-law tails in the distribution of $q$ that can be a confounding factor. If the tail exponent is equal to $\varpi$ (see section 4.4.1), one expects a crossover value $a_{c}(n)$ given by $\varpi/2n$, i.e. when $\mathbb{E}[q^{2na}]$ diverges. Although such a mechanism may certainly play a role, it comes in parallel with the dependence of $\mu_{q}$ on $q$. We will compare these predictions with empirical data in the next section. Although the model is over-simplified, we will see that it captures the data semi-quantitatively. Typically, $a_{c}(1)$ is found to be around $2$. With $\mu_{m}=3/2$, we find $\lambda\sigma_{\ell}^{2}\approx 1/8$, an estimate that will match other observations.